Any help on this question would be greatly appreciated. I need to give my answer in terms of the radius of earth. Thanks!

The period of a pendulum oscillating on the surfac
Answer #1

Period = T = 2pi*sqrt(L/g)

Tp/Te = sqrt(ge/gp) =======> 1.4 = sqrt(9.8/gp) ====> gp = 5 m/s^2

Now radius = sqrt[GM/gp] = sqrt[(6.6726*10-11*0.6*5.972*10^24)/5] = 6915093.7 m = 6.92*10^6 m